Katie likes to wear odd socks. She can choose two non-matching socks from a drawer with six different, separated pairs of socks in how many ways?
Answer by jim_thompson5910(35256)

You have 6*2 = 12 socks total.

Let's say you have two slots: AB
Slot A represents the first sock, B is the second sock.

You have 12 choices for slot A, and 10 choices for slot B. Why 10? Because once you pick that first sock, you go from 12 to 12-1 = 11 socks...BUT...you can't pick the sock that matches with the sock for slot A. So you take that matching sock out to get 11-1 = 10 choices for slot B.

Multiply the choices out: 12*10 = 120


There are 120 ways to pick 2 socks that are non-matching pairs.